OpenStudy (anonymous):

if f(x)=cosx and g(x)=9^x, what is the range of g(f(x))?

OpenStudy (anonymous):

\[\large g(f(x))=9^{\cos x}\] \(\cos x\) is bounded between 1 and -1, which means \(f(g(x))\) is bounded between \(9^1\) and \(9^{-1}\).

OpenStudy (jhannybean):

\(9^{-1} = \cfrac{1}{9}\)

OpenStudy (jhannybean):

\[\large \cfrac{1}{9} \le 9^{cos(x)} \le 9\]

OpenStudy (anonymous):

OOOOH !....

OpenStudy (jhannybean):

:)

OpenStudy (anonymous):

thanks..

OpenStudy (anonymous):

:)

OpenStudy (jhannybean):

No problemo